An object of 6 kg mass was at rest. After that a 30 N force is imposed on it. What will be
kinetic energy after 10 sec time?
1
Expert's answer
2012-08-10T09:52:01-0400
Let's find the acceleration of an object using the second Newton's law:

F = ma ==> a = F/m = 30N /6kg = 5 m/s².

The velocity of an object after 10 seconds will be

V = at = 5m/s² * 10s = 50m/s.

So, the kinetic energy of an object after 10 seconds will be

E = mV²/2 = 6kg * (50m/s)² /2 = 7500J.